OpenStudy (anonymous):

the expression is a perfect square if b>0 find b y=x^2+bx+49

OpenStudy (anonymous):

14

OpenStudy (anonymous):

\[(a+c)^2=a^2+2ac+c^2\] here \[a=x,c=7\] so \[2c=14\]